# Andreev spin qubit protected by Franck-Condon blockade

### v1

@@ Abstract
Andreev levels localized in a weak link between two superconductors can trap a superconducting quasiparticle. If there is a spin-orbit coupling in the link, the spin of the quasiparticle couples to the Josephson current. This effect can be leveraged to control and readout the spin of the quasiparticle thus using it as a qubit. One of the factors limiting the performance of such an Andreev spin qubit is spin relaxation. Here, we theoretically demonstrate that the relaxation lifetime can be enhanced by utilizing the coupling between the Andreev spin and the supercurrent in a transmon circuit. The coupling ensures that the flip of the quasiparticle spin can only happen if it is accompanied by the excitation of multiple plasmons, as dictated by the Frank-Condon principle. This blocks spin relaxation at temperatures small compared to plasmon energy.
Introduction. The spin of a single quasiparticle trapped in a Josephson junction can couple to the supercurrent via the spin-orbit interaction [1]. Following several years of theoretical work [1, 2, 3, 4, 5, 6, 7], this effect has been experimentally established via microwave spectroscopy measurements [8, 9]. It is a key element for the operation of Andreev spin qubits [10, 11, 12]. These novel qubits have multiple properties favorable for quantum computing: a $\mu\mathrm{m}$-scale spatial footprint, large relative anharmonicity, and a direct interaction with microwave fields mediated by the spin-supercurrent coupling. The latter coupling also allows for cm-scale qubit connectivity [12] which is useful for quantum simulations [13].
In the current experimental realizations of Andreev spin qubits, however, the benefits listed above are outweighed by short decoherence times [11, 14, 15]. Both dephasing and relaxation times of Andreev spin qubits are much shorter than those of state-of-the-art semiconductor spin qubits [16, 17]. At least in part, these times could be improved by updating the material platform used in Andreev spin qubit devices. So far, all measured devices were based on InAs/Al nanowires, which are far from optimal for spin coherence [16]. Material considerations aside, one may wonder: in the spirit of hardware-protected qubits [18], is it possible to enhance the coherence of the Andreev spin qubit by circuit design?
@@
In this Letter, we show that the relaxation time of the Andreev spin qubit can be enhanced by shunting it with a capacitor, as in a transmon circuit [19]. Then, provided the spin-dependent supercurrent is large enough, the transmon ground state wave-functions for opposite spin orientations become disjoint, separated by a distance $\sim 2\phi_{0}$ larger than their extent $\sim 2\phi_{c}$, see Fig. 1. In this situation, the matrix elements for spin-flip transitions are exponentially suppressed with $(\phi_{0}/\phi_{c})^{2}$, leading to an enhancement of the spin relaxation time. The suppression of transition matrix elements is governed by Franck-Condon principle [20], and we therefore refer to it as the Franck-Condon blockade, by analogy with a related phenomenon in molecular transport [21, 22]. The pure dephasing time of the Andreev spin qubit is neither improved nor worsened by Frank-Condon blockade.
We build a systematic theory of Frank-Condon blockade of Andreev spin qubit relaxation. We describe how the protection can be compromised by finite temperature. In fact, we show that the residual spin relaxation comes from a thermally-activated process where a spin-flip is accompanied by an excitation of several plasmonic quanta. Despite requiring energy, this process has a larger matrix element compared to the excitation-less process, as described by the Frank-Condon principle. Additionally, we identify experimentally observable consequences of the Frank-Condon blockaded regime: the appearance of plasmonic transitions in response to an external drive acting on the spin and a staircase-shaped dependence of spin lifetime on magnetic field. Finally, we provide a guideline for future experiments by discussing the device design of the protected Andreev spin qubit. Throughout the paper we use units with $k_{B}=\hbar=1$.
Model and main idea. We consider a circuit in which the junction trapping the spin-$1/2$ quasiparticle is shunted by capacitance $C$, see Fig. 1. Adopting a minimal model for the Josephson energy [5], the Hamiltonian of the circuit is given by
$H=4E_{c}(i\partial_{\phi}+n_{g})^{2}+E_{0}\cos(\phi)+E_{\textrm{so}}\sigma_{z}\sin(\phi)\,.$ (1)
@@
The last term of Eq. (1) arises due to spin-orbit interaction. It is the simplest coupling between spin and superconducting phase compatible with time-reversal symmetry (which sends $\phi\to-\phi$ and $\sigma_{z}\to-\sigma_{z}$) and the compactness of $\phi$. In principle, the direction $\hat{n}$ of the spin-orbit field may depend on the phase difference, so that $\sigma_{z}$ should be replaced by a more general operator $\vec{\sigma}\cdot\hat{n}(\phi)$ with $\hat{n}(\phi)=\hat{n}(-\phi)$. This replacement opens new channels of spin relaxation, as it couples the spin to charge fluctuations via $n_{g}$. This, however, does not qualitatively affect our results regarding the Frank-Condon blockade. We therefore resort to the simple form of the coupling Hamiltonian given by Eq. (1).
@@
To describe Frank-Condon blockade, we introduce a parameter $\xi_{0}=(\phi_{0}/\phi_{c})^{2}$. If $\xi_{0}~{}\gg~{}1$, the lowest-level wave functions with opposite spin are disjoint. Their overlap is exponentially small in $\xi_{0}$, since it is the integral of a product of two Gaussian wave functions of width $\sim\phi_{c}$ at a distance $\sim\phi_{0}$, see Fig. 1. We call regime $\xi_{0}\gg 1$ “the strong coupling regime”, referring to the strength of coupling between the spin and the supercurrent. To achieve the strong coupling regime it is sufficient to fulfill two conditions: (i) $\omega_{p}\gg E_{c}$ such that the circuit is in the transmon regime [19] with small zero-point phase fluctuations ($\phi_{c}\ll 1$), (ii) $E_{\textrm{so}}\gtrsim E_{0}$ such that the separation between potential minima for opposite spin is large ($\phi_{0}\sim 1$).
@@
Figure 2: (a): Temperature dependence of the spin-flip rate at $J_{z}=0$. We assume that the system is initialized in the plasmon ground state $k=0$ for one of the two spins. The solid line is Eq. (6), the purple dashed line is the approximate expression (8), and the horizontal dashed line is the $T=0$ result (5). The parameters are $E_{c}/E_{\textrm{so}}=0.035$ and $E_{0}/E_{\textrm{so}}=1$, yielding $\omega_{p}/E_{\textrm{so}}\approx 0.63$ and $\xi_{0}\approx 2.8$. (b)-(c): Dependence of the spin-flip rate on temperature and magnetic field parallel to the spin-orbit axis, starting from either spin down (b) or spin up (c). The rates display step-like features as a function of the magnetic field, which are washed out by increasing temperature. The vertical and horizontal lines correspond to the line cuts shown in panels (a) and (d) respectively. (d): Magnetic field dependence of the spin-flip rate for initial states with opposite spin (initialized in a state with $k=0$). The solid curves are computed at $T/\omega_{p}=0.1$, while the dashed ones at $T/\omega_{p}=1$. In the low temperature curves, the steps occur at half-integer values of $J_{z}/\omega_{p}$, and have heights given by the coefficients $w_{k}$. The results presented in this figure are derived for a simple bath of two-level systems, with a constant density of states and energy-independent coupling to the Andreev spin.
@@
Assume now that the Andreev spin is coupled to a vector field $\vec{J}$, i.e., we add an extra term $H_{J}=\vec{J}\cdot\vec{\sigma}$ to the Hamiltonian in Eq. (1). This vector field can describe either an intentionally applied magnetic field or the uncontrolled noise stemming from the environment of the spin. To demonstrate the essence of the Frank-Condon blockade, we can momentarily leave the origin of $\vec{J}$ unspecified. If we project $H_{J}$ onto the $k=0$ subspace, we obtain the effective Hamiltonian:
$H_{\textrm{eff}}=e^{-\xi_{0}/2}(J_{x}\sigma_{x}+J_{y}\sigma_{y})+J_{z}\sigma_{z}\,.$ (4)
@@
We see that the matrix elements of $H_{J}$ between ground states of opposite spins are suppressed a factor $e^{-\xi_{0}/2}$ which is small when $\xi_{0}\gg 1$. Thus, the external field is not effective in flipping the Andreev spin, and the qubit encoded in this spin is protected from relaxation. This is the main idea of our proposal.
@@
Before assessing the conditions for the validity of Eq. (4), let us show explicitly that it leads to the enhancement of spin relaxation. To this end, we assume that the field ${\vec{J}}$ encodes a spin-flipping environment of the Andreev spin. Realistically, such environment will be constituted by nuclear spins in combination with phonons [24], exchanged-coupled itinerant quasiparticles in the superconducting leads [25], magnetic field noise, and charge noise. However, on a qualitative level, the nature of the environment is not important for the discussion of Frank-Condon blockade. Therefore, to be concrete, we adopt a simplified model in which the environment is a collection of two-level systems of an unspecified origin. We assume that this bath has a constant density of states $\nu_{0}$ and each two-level system couples to $\sigma_{x}$ with the same strength $g$. Using Fermi’s golden rule, we then find
@@
Temperature dependence of spin relaxation rate. The effective Hamiltonian given by Eq. (4) does not take into account the plasmonic excitations (states with $k>0$). At the same time, the full version of the coupling Hamiltonian Hamiltonian $H_{J}$ mixes the low-energy states, $k=0$, with the excited states, $k>0$. For this reason, when describing the dynamics of the Andreev spin, neglecting the excited states is only allowed when the relevant energy scales (such as the temperature or $\vec{J}$) are much smaller than $\omega_{p}$. Otherwise, a more careful calculation of the spin-flip dynamics is warranted. In fact, as we now demonstrate, a temperature as low as $\omega_{p}/\ln(\xi_{0})$ can significantly compromise the protection.
To reveal the role of finite temperature, we resort to the two-level system model for the bath described above. For simplicity, we assume that the system is initialized in the plasmon ground state, $k=0$, for one of the spin directions. The total spin-flip rate can then be evaluated as
@@
According to this equation, Eq. (5) is accurate as long as the temperature satisfies $T\ll\omega_{p}/\ln\xi_{0}$. Otherwise, Eq. (5) underestimates the transition rate. The rate is enhanced due to the possibility of a spin-flip with a simultaneous excitation of $k$ plasmons. Even though this process is penalized by a factor $e^{-\beta\omega_{p}k}$, it has a much larger matrix element compared to that unaccompanied by plasmon excitations. In the regime $T\gg\omega_{p}$, we obtain $\Gamma_{\uparrow\rightarrow\downarrow}=\Gamma_{0}$ since $e^{\beta\omega_{p}k}\approx 1$ for all $k$ and the following sum rule holds:
@@
Effect of a static magnetic field. Next, we discuss experimental signatures of a protected regime of the Andreev spin qubit for $T\ll\omega_{p}$. The first such signature is the unusual magnetic field dependence of the spin-flip rate. To see this dependence, we assume that a static magnetic field is applied parallel to the spin-orbit axis and so we include a term $J_{z}\sigma_{z}$ in the Hamiltonian. Then, Eq. (6) generalizes to
@@
[…]
The Franck-Condon factors of Eq. (7) also determine the splitting of Kramers-degenerate excited states due to magnetic field perpendicular to the spin-orbit axis. Indeed, according to first-order perturbation theory, term $H_{J}=J_{x}\sigma_{x}$ in the Hamiltonian splits the energy of the $k$-th level by $\delta E_{k}\approx J_{x}g_{kk}=J_{x}e^{-\xi_{0}/2}L_{k}(\xi_{0})$, with $L_{k}$ the $k$-th Laguerre polynomial.
[…]

### v2

@@ Abstract
Andreev levels localized in a weak link between two superconductors can trap a superconducting quasiparticle. If there is a spin-orbit coupling in the link, the spin of the quasiparticle couples to the Josephson current. This effect can be leveraged to control and readout the spin of the quasiparticle thus using it as a qubit. One of the factors limiting the performance of such an Andreev spin qubit is spin relaxation. Here, we theoretically demonstrate that the relaxation lifetime can be enhanced by utilizing the coupling between the Andreev spin and the supercurrent in a transmon circuit. The coupling ensures that the flip of the quasiparticle spin can only happen if it is accompanied by the excitation of multiple plasmons, as dictated by the Franck-Condon principle. This blocks spin relaxation at temperatures small compared to plasmon energy.
I Introduction
The spin of a single quasiparticle trapped in a Josephson junction can couple to the supercurrent via the spin-orbit interaction [1]. Following several years of theoretical work [1, 2, 3, 4, 5, 6, 7], this effect has been experimentally established via microwave spectroscopy measurements [8, 9]. It is a key element for the operation of Andreev spin qubits [10, 11, 12]. These novel qubits have multiple properties favorable for quantum computing: a $\mu\mathrm{m}$-scale spatial footprint, large relative anharmonicity, and a direct interaction with microwave fields mediated by the spin-supercurrent coupling. The latter coupling also allows for cm-scale qubit connectivity [12] which is useful for quantum simulations [13].
In the current experimental realizations of Andreev spin qubits, however, the benefits listed above are outweighed by short decoherence times [11, 14, 15]. Both dephasing and relaxation times of Andreev spin qubits are much shorter than those of state-of-the-art semiconductor spin qubits [16, 17]. So far, all measured devices were based on InAs/Al nanowires, which are far from optimal for spin coherence [16], in large part due to the presence of a nuclear spin bath.
A strategy to increase the dephasing time is to update the material platform used in Andreev spin qubit devices to one with less nuclear spins. A promising example is isotopically purified germanium [18]. Notably, Ge-based quantum dot Josephson junctions have been realized [19], and theoretical proposals for Andreev spin qubits in the material have been developed [20].
In contrast, the origin of relaxation in Andreev spin qubits is still under active investigation [14], and therefore a strategy to improve relaxation time is still lacking. Material considerations aside, one may wonder: in the spirit of hardware-protected qubits [21], is it possible to enhance the relaxation time of the Andreev spin qubit by circuit design?
@@ I Introduction
In this work, we show that the relaxation time of the Andreev spin qubit can be enhanced by shunting it with a capacitor, as in a transmon circuit [22]. Then, provided the spin-dependent supercurrent is large enough, the transmon ground state wave-functions for opposite spin orientations become disjoint, separated by a distance $\sim 2\phi_{0}$ larger than their extent $\sim 2\phi_{c}$, see Fig. 1. In this situation, the matrix elements for spin-flip transitions are exponentially suppressed with $(\phi_{0}/\phi_{c})^{2}$, leading to an enhancement of the spin relaxation time. The suppression of transition matrix elements is governed by Franck-Condon principle [23], and we therefore refer to it as the Franck-Condon blockade, by analogy with a related phenomenon in molecular transport [24, 25]. The pure dephasing time of the Andreev spin qubit is neither improved nor worsened by Franck-Condon blockade.
We build a systematic theory of Franck-Condon blockade of Andreev spin qubit relaxation. We describe how the protection can be compromised by finite temperature. In fact, we show that the residual spin relaxation comes from a thermally-activated process where a spin-flip is accompanied by an excitation of several plasmonic quanta. Despite requiring energy, this process has a larger matrix element compared to the excitation-less process, as described by the Franck-Condon principle. Additionally, we identify experimentally observable consequences of the Franck-Condon blockaded regime: the appearance of plasmonic transitions in response to an external drive acting on the spin and a staircase-shaped dependence of spin lifetime on magnetic field. Finally, we provide a guideline for future experiments by discussing the device design of the protected Andreev spin qubit. Throughout the paper we use units with $k_{B}=\hbar=1$.
II Model and main idea
We consider a circuit in which the junction trapping the spin-$1/2$ quasiparticle is shunted by capacitance $C$, see Fig. 1. Adopting a minimal model for the Josephson energy [5], the Hamiltonian of the circuit is given by
$H_{0}=4E_{c}(i\partial_{\phi}+n_{g})^{2}+E_{0}\cos(\phi)+E_{\textrm{so}}\sigma_{z}\sin(\phi)\,.$ (1)
@@ II Model and main idea
The last term of Eq. (1) arises due to spin-orbit interaction. It is the simplest coupling between spin and superconducting phase compatible with time-reversal symmetry (which sends $\phi\to-\phi$ and $\sigma_{z}\to-\sigma_{z}$) and the compactness of $\phi$. In principle, the direction $\hat{n}$ of the spin-orbit field may depend on the phase difference, so that $\sigma_{z}$ should be replaced by a more general operator $\vec{\sigma}\cdot\hat{n}(\phi)$ with $\hat{n}(\phi)=\hat{n}(-\phi)$. This replacement opens new channels of spin relaxation, as it couples the spin to charge fluctuations via $n_{g}$. This, however, does not qualitatively affect our results regarding the Franck-Condon blockade. We therefore resort to the simple form of the coupling Hamiltonian given by Eq. (1).
@@
To describe Franck-Condon blockade, we introduce a parameter $\xi_{0}=(\phi_{0}/\phi_{c})^{2}$. If $\xi_{0}~\gg~1$, the lowest-level wave functions with opposite spin are disjoint. Their overlap is exponentially small in $\xi_{0}$, since it is the integral of a product of two Gaussian wave functions of width $\sim\phi_{c}$ at a distance $\sim\phi_{0}$, see Fig. 1. We call regime $\xi_{0}\gg 1$ “the strong coupling regime”, referring to the strength of coupling between the spin and the supercurrent. To achieve the strong coupling regime it is sufficient to fulfill two conditions: (i) $\omega_{p}\gg E_{c}$ such that the circuit is in the transmon regime [22] with small zero-point phase fluctuations ($\phi_{c}\ll 1$), (ii) $E_{\textrm{so}}\gtrsim E_{0}$ such that the separation between potential minima for opposite spin is large ($\phi_{0}\sim 1$).
@@
Figure 2: (a): Temperature dependence of the spin-flip rate at $J_{z}=0$. We assume that the system is initialized in the plasmon ground state $k=0$ for one of the two spins. The solid line is Eq. (10), the purple dashed line is the approximate expression (12), and the horizontal dashed line is the $T=0$ result (9). The parameters are $E_{c}/E_{\textrm{so}}=0.035$ and $E_{0}/E_{\textrm{so}}=1$, yielding $\omega_{p}/E_{\textrm{so}}\approx 0.63$ and, using Eq. (2) and (3), $\xi_{0}\approx 2.8$. (b)-(c): Dependence of the spin-flip rate on temperature and magnetic field parallel to the spin-orbit axis, starting from either spin down (b) or spin up (c). The rates display step-like features as a function of the magnetic field, which are washed out by increasing temperature. The vertical and horizontal lines correspond to the line cuts shown in panels (a) and (d) respectively. (d): Magnetic field dependence of the spin-flip rate for initial states with opposite spin (initialized in a state with $k=0$). The solid curves are computed at $T/\omega_{p}=0.1$, while the dashed ones at $T/\omega_{p}=1$. In the low temperature curves, the steps occur at half-integer values of $J_{z}/\omega_{p}$, and have heights given by the coefficients $w_{k}$. The results presented in this figure are derived for a simple bath of two-level systems, with a constant density of states and energy-independent coupling to the Andreev spin.
@@
Assume now that the Andreev spin is coupled to a vector field $\vec{J}$. This vector field can describe either an intentionally applied magnetic field or the uncontrolled noise stemming from the environment of the spin. To demonstrate the essence of the Franck-Condon blockade, we can momentarily leave the origin of $\vec{J}$ unspecified.
The coupling between the Andreev spin qubit and $\vec{J}$ can be incorporated into the model by adding an extra term
$H_{J}=\vec{J}\cdot\vec{\sigma}$ (4)
to the Hamiltonian $H_{0}$ of Eq. (1). Expressed in the basis of eigenstates of $H_{0}$, labeled by the discrete quantum number $k$ and the spin quantum number $\sigma=\uparrow,\downarrow$, the total Hamiltonian $H=H_{0}+H_{J}$ reads
$H$ $=\sum_{k,\sigma}E_{k}\,\left|k,\sigma\right\rangle\left\langle k,\sigma\right|$
$+\vec{J}\,\cdot\,\sum_{kk^{\prime}}\sum_{\sigma,\sigma^{\prime}}\,\left\langle k,\sigma\right|\vec{\sigma}\left|k^{\prime},\sigma^{\prime}\right\rangle\left|k,\sigma\right\rangle\left\langle k^{\prime},\sigma^{\prime}\right|\,.$ (5)
We emphasize again that the unperturbed energy levels $E_{k}$ are Kramers degenerate and have wave functions centered around $\pi\pm\phi_{0}$, depending on $\sigma$. This relative shift decreases the overlap of wave functions with opposite spin $\sigma$ and therefore suppresses the matrix elements $\left\langle k,\sigma\right|\vec{\sigma}\left|k^{\prime},\sigma^{\prime}\right\rangle$ if $\sigma\neq\sigma^{\prime}$.
As mentioned above, in the harmonic limit $\omega_{p}\gg E_{c}$, the quantum number $k$ identifies the number of plasmon excitations and the wave functions are those of the harmonic oscillator. In this limit, if we project $H$ onto the Kramers-degenerate ground state space, that is onto the $k=0$ subspace, we obtain the effective Andreev spin Hamiltonian
$H_{\textrm{eff}}=E_{0}+e^{-\xi_{0}/2}(J_{x}\sigma_{x}+J_{y}\sigma_{y})+J_{z}\sigma_{z}\,.$ (6)
@@
We see that the matrix elements between ground states of opposite spins are suppressed a factor $e^{-\xi_{0}/2}$ which is small when $\xi_{0}\gg 1$. Thus, the external field is not effective in flipping the Andreev spin, and the qubit encoded in this spin is protected from relaxation. This is the main idea of our proposal.
III Spin relaxation rate
@@
We now show that Eq. (6) leads to the suppression of spin relaxation. To this end, we assume that the field $\vec{J}$} encodes a spin-flipping environment of the Andreev spin. Realistically, such environment will be constituted by nuclear spins in combination with phonons [27], exchanged-coupled itinerant quasiparticles in the superconducting leads [28], magnetic field noise, and charge noise. However, on a qualitative level, the nature of the environment is not important for the discussion of Franck-Condon blockade. Therefore, to be concrete, we adopt a simplified model in which the environment is a collection of two-level systems of an unspecified origin.
The Hamiltonian of this simplified model is
$H_{\textrm{tot}}$ $=H_{0}+\sum_{i}\,\epsilon_{i}\sigma_{z,i}+\sigma_{x}\,J_{x}\,$, (7)
$J_{x}$ $=g\sum_{i}\sigma_{x,i}\,$, (8)
where $H_{0}$ is the Andreev spin qubit Hamiltonian of Eq. (1), $i$ runs over all the two level systems in the environment, and $\epsilon_{i}$ is the energy associated with each two level system. We assume for simplicity that the energies are uniformly distributed, so that the environment has a constant density of states $\nu_{0}$. The last term describes the coupling between the Andreev spin and the two level systems, and has been chosen in a form that allows Andreev spin flips induced by the environment. Again for simplicity, we have assumed that the coupling constant $g$ in the last term does not depend on $i$. These simplifying assumptions can be relaxed without qualitative consequences for the results that follow.
We can compute the spin relaxation rate $\Gamma_{\uparrow\rightarrow\downarrow}$ as the transition rate between the two degenerate ground states of $H_{0}$. Within this simple bath model, using Fermi’s golden rule, we find
@@
III.1 Temperature dependence
The effective Hamiltonian given by Eq. (6) does not take into account the plasmonic excitations, that is all the states with $k>0$. At the same time, the full version of the coupling Hamiltonian $H_{J}$ mixes the low-energy states, $k=0$, with the excited states, $k>0$. For this reason, when describing the dynamics of the Andreev spin, neglecting the excited states is only allowed when the temperature is smaller than $\omega_{p}$. Otherwise, a more careful calculation of the spin-flip dynamics is warranted. In fact, as we now demonstrate, a temperature as low as $\omega_{p}/\ln(\xi_{0})$ can significantly compromise the protection.
To reveal the role of finite temperature, we resort to the same two-level system model for the bath described above. For simplicity, we assume that the system is initialized in the plasmon ground state, $k=0$, for one of the spin directions. The total spin-flip rate can then be evaluated as
@@ III.1 Temperature dependence
According to this equation, Eq. (9) is accurate as long as the temperature satisfies $T\ll\omega_{p}/\ln\xi_{0}$. Otherwise, Eq. (9) underestimates the transition rate. The rate is enhanced due to the possibility of a spin-flip with a simultaneous excitation of $k$ plasmons. Even though this process is penalized by a factor $e^{-\beta\omega_{p}k}$, it has a much larger matrix element compared to that for a process unaccompanied by plasmon excitations. In the regime $T\gg\omega_{p}$, we obtain $\Gamma_{\uparrow\rightarrow\downarrow}=\Gamma_{0}$ since $e^{\beta\omega_{p}k}\approx 1$ for all $k$ and the following sum rule holds:
@@
The temperature dependence of the spin relaxation rate is illustrated in Fig. 2a, which includes a comparison between Eq. (10) and its approximation Eq. (12).
@@
IV Experimental signatures
We now discuss two experimental signatures of a protected regime of the Andreev spin qubit for $T\ll\omega_{p}$. The first such signature is the unusual magnetic field dependence of the spin-flip rate. Another is the appearance of plasmonic branches in the spectrum of a driven Andreev spin qubit.
IV.1 Effect of a static magnetic field
To see this dependence, we assume that a static magnetic field is applied (anti-)parallel to the spin-orbit axis and so we include a term $J_{z}\sigma_{z}$ in the Hamiltonian. We note that the magnitude and sign of $J_{z}$ are determined by the external magnetic field through the $g$-factor of the Andreev spin, which may depend on the material platform and the geometry of the junction [11, 19]. Then, Eq. (10) generalizes to
@@ IV.1 Effect of a static magnetic field
From this equation, we see that for $J_{z}<0$ and $T\lesssim|J_{z}|$, the spin-flip rate is additionally suppressed compared to Eq. (9). Indeed, an extra energy of $2J_{z}$ is required to excite the initial spin-up state into a spin-down state, even without exciting plasmons. For positive $J_{z}$, instead, an increasing magnetic field sequentially lifts the energy of the $\left|0\uparrow\right\rangle$ state above that of the $\left|k\downarrow\right\rangle$ state. Each time this happens, a new relaxation channel opens up for the spin-up state. Hence, the spin-flip rate increases in steps [29] centered around the points $J_{z}=k\omega_{p}/2$, see Fig. 2d. The $k$-th step has height $\Gamma_{0}w_{k}$. The relaxation rate with an opposite initial spin has an inverted field dependence, $\Gamma_{\downarrow\to\uparrow}(J_{z})=\Gamma_{\uparrow\to\downarrow}(-J_{z})$, as illustrated in Fig. 2b-c.
The Franck-Condon factors of Eq. (11) also determine the splitting of Kramers-degenerate excited states due to magnetic field perpendicular to the spin-orbit axis. Indeed, according to first-order perturbation theory, the term $H_{J}=J_{x}\sigma_{x}$ in the Hamiltonian splits the energy of the $k$-th level by $\delta E_{k}\approx J_{x}g_{kk}=J_{x}e^{-\xi_{0}/2}L_{k}(\xi_{0})$, with $L_{k}$ the $k$-th Laguerre polynomial.
IV.2 Driven excitation of plasmonic branches
Consider a driving field coupled to $\sigma_{x}$, $H_{J}=J_{x}\cos(\omega t)\sigma_{x}$. If the frequency of the drive matches $2J_{z}$ then Rabi oscillations develop between states $|0\uparrow\rangle$ and $|0\downarrow\rangle$. However, $2J_{z}$ is not the only frequency activating the spin-flip process. Namely, if the driving frequency is resonant with a plasmonic transition, $\omega=k\omega_{p}+2J_{z}$, the driving field will selectively induce coherent oscillations between the states $\left|0\uparrow\right\rangle$ and $\left|k\downarrow\right\rangle$, with a Rabi frequency